# Consistent3D: Towards Consistent High-Fidelity Text-to-3D Generation with Deterministic Sampling Prior

### v1

@@ Abstract
Score distillation sampling (SDS) and its variants have greatly boosted the development of text-to-3D generation, but are vulnerable to geometry collapse and poor textures yet. To solve this issue, we first deeply analyze the SDS and find that its distillation sampling process indeed corresponds to the trajectory sampling of a stochastic differential equation (SDE): SDS samples along an SDE trajectory to yield a less noisy sample which then serves as a guidance to optimize a 3D model. However, the randomness in SDE sampling often leads to a diverse and unpredictable sample which is not always less noisy, and thus is not a consistently correct guidance, explaining the vulnerability of SDS. Since for any SDE, there always exists an ordinary differential equation (ODE) whose trajectory sampling can deterministically and consistently converge to the desired target point as the SDE, we propose a novel and effective “Consistent3D” method that explores the ODE deterministic sampling prior for text-to-3D generation. Specifically, at each training iteration, given a rendered image by a 3D model, we first estimate its desired 3D score function by a pre-trained 2D diffusion model, and build an ODE for trajectory sampling. Next, we design a consistency distillation sampling loss which samples along the ODE trajectory to generate two adjacent samples and uses the less noisy sample to guide another more noisy one for distilling the deterministic prior into the 3D model. Experimental results show the efficacy of our Consistent3D in generating high-fidelity and diverse 3D objects and large-scale scenes, as shown in Fig. 1. The codes are available at https://github.com/sail-sg/Consistent3D.
@@
footnotetext: $*{}^{*}$orresponding author.
@@ 1 Introduction
To address this critical issue, in this paper, we propose a novel and effective method, dubbed“Consistent3D”, which guides text-to-3D generation using deterministic sampling prior and can generate high-fidelity 3D objects and large-scale scenes. As illustrated in Fig. 2, for these unpredictable and uncontrollable trajectories sampled from SDE solution distribution, there theoretically always exists a corresponding ordinary differential equation (ODE) whose trajectory shares the same marginal distributions [41] with the SDE solution. Importantly, the ODE trajectory is deterministic and consistently converges to the same target point as the SDE. Sampling along the ODE trajectory guarantees a predictable and deterministic next point, which always directs towards the desired target and can provide a reliable and consistent guidance. This motivated us to explore the text-to-3D generation from the ODE deterministic sampling perspective.
@@ 2 Related Works
Text-to-3D Generation stands for generating 3D contents from a given text description. Current 3D generative models [29, 15], usually work in a single object category and suffer from limited diversity due to the lack of large-scale 3D datasets. To achieve open-vocabulary 3D generation, pioneered by DreamFusion [31], several approaches propose to lift text-image diffusion models [34] for 3D generation [54, 47, 55]. The key mechanism of such approaches is the score distillation sampling (SDS), where diffusion priors are used to supervise the optimization of a 3D representation. The following works continue to further improve the stability and fidelity of generation of various aspects, e.g., advanced 3D representation [3, 43, 51, 46], coarse-to-fine training strategy [20, 55, 47] and 3D-aware diffusion priors [54, 38, 19, 22, 44].
@@ 4 Method
Here we elaborate on our proposed Consistent3D for effective text-to-3D generation. In Sec. 4.1, we reveal the underlying mechanism of Score Distillation Sampling (SDS) which aims to approximate the SDE sampling process and motivates our proposed methods. Then in Sec. 4.2, we introduce Consistency Distillation Sampling (CDS), a loss designed to efficiently distill deterministic sampling priors for text-to-3D generation, accompanied with a theoretical justification regarding the error bound. Finally, we present how to use our CDS to build our text-to-3D generation framework, termed as “Consistent3D”, in Sec. 4.3.
@@
       $t_{2}\leftarrow t_{\max}-(t_{\max}-t_{\min})\sqrt{{i}/{N}}$ Sample camera pose $\pi$ Sample $t_{1}\in{\mathcal{U}}[t_{2}+\delta,t_{2}+\Delta]$ $\mathbf{x}_{\pi}\leftarrow g(\bm{\theta},\pi)$ $\mathbf{x}_{t_{1}}\leftarrow\mathbf{x}_{\pi}+\sigma_{t_{1}}\bm{\epsilon}^{*}$ $\mathbf{d}_{i}\leftarrow\left(\mathbf{x}_{t_{1}}-D_{\phi}(\mathbf{x}_{t_{1}},t_{1},y)\right)/\sigma_{t_{1}}$ $\hat{\mathbf{x}}_{t_{2}}\leftarrow\mathbf{x}_{t_{1}}+(\sigma_{t_{2}}-\sigma_{t_{1}})\mathbf{d}_{i}$ $\hat{\mathbf{x}}_{0}\leftarrow\mathbf{x}_{\pi}+\operatorname{sg}\left[\sigma_{t_{1}}(\bm{\epsilon}^{*}-\mathbf{d}_{i})\right]$ ${\mathcal{L}}_{\text{CDS}}(\bm{\theta};\pi)\leftarrow\lambda(t_{2})\lVert\hat{\mathbf{x}}_{0}-\operatorname{sg}\left[D_{\phi}(\hat{\mathbf{x}}_{t_{2}},t_{2},y)\right]\rVert_{2}^{2}$ $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}{\mathcal{L}}_{\text{CDS}}(\bm{\theta};\pi)$
@@ 4.1 Revisit Score Distillation Sampling
Firstly, we discretize the reverse SDE in Eq. (2) and perform the stochastic sampling process following Ho et al. [11], which results in the SDE solution trajectory defined as
@@ 4.2 Consistency Distillation Sampling
Specifically, we define the time step schedule of $t_{2}$ according to the current training iteration:
@@
[…]
@@ 4.3 Consistent3D
NeRF Optimization Stage. We adopt multi-resolution hash grids, Instant NGP [28] to parameterize the scene by density and color with MLPs, which improves training and rendering efficiency. We follow Magic3D on density bias initialization, camera and light augmentation. In addition, we use orientation loss [31] and 2D normal smooth loss [24]. At this stage, we render $64\times 64$ images and use our proposed CDS as guidance. We set $t_{\max}=0.98$, $t_{\min}=0.1$, $\delta=0.1$, and $\Delta=0.2$.
Mesh Refinement Stage. We convert the neural field into Signed Distance Field (SDF) by subtracting it with a fixed threshold and then optimizing a high-resolution DMTet [37]. We also initialize the volume texture field directly with the color field from the coarse stage. In addition, we use normal consistency loss and Laplacian smoothness loss. In the refinement stage, we render $512\times 512$ images and set $t_{\max}=0.5$, $t_{\min}=0.02$, $\delta=0.1$, and $\Delta=0.1$. Unlike previous works [47, 3], our approach is capable of refining both geometry and texture simultaneously for efficiency, thanks to the consistent guidance provided by our CDS.
@@ 5.1 Implementation Details
Consistent3D is implemented in PyTorch with a single NVIDIA A100 GPU based on threestudio [9]. We use the Adan [49] optimizer with a learning rate of $0.05$ for grid encoder and $0.005$ for other parameters. Additionally, we implemented a weight decay of $2\times 10^{-8}$. Further implementation details are provided in Sec. B.1.
@@ 5.3 Comparison with the State-of-The-Art
Quantitative Results. In Table 1, we report the results of CLIP R-Precision [32] for 3D objects generated using $40$ randomly selected text prompts from the DreamFusion gallery. See more details in Sec. B.1. Each 3D object is rendered from $120$ viewpoints with a uniform azimuth angle. The CLIP R-Precision score is computed by averaging the similarity scores between each rendered view and the corresponding text prompt. Our quantitative analysis demonstrates the superior performance of our method.
@@ 5.4 Ablation Study
We present an ablation study to evaluate the effects of various components in our approach. We conduct experiments with the following configurations: (a) a random time step schedule in DreamFusion [31]; (b) a predetermined time step schedule from Consistency Distillation [42]; (c) varied random noise in each iteration; and (d) our proposed method incorporating all components, with results summarized in Fig. 6. The results in Fig. 6(a) reveal that a random time step schedule detrimentally affects both geometry and texture modeling, since it disrupts established rules of sampling process. Fig. 6(b) suggests that a predetermined time step schedule is suboptimal for optimization-based methods, since gradient descent does not ensure monotonic optimization progress. This implies that minor randomness helps to accommodate these variations. Fig. 6(c) shows that fixed noise aids in better convergence by providing a consistent perturbation in each iteration.
@@ 6 Conclusion
In this work, we first connect Score Distillation Sampling (SDS), a leading text-to-3D generation approach, with the solution trajectory sampling of a stochastic differential equation (SDE). This connection helps us to understand the vulnerability in SDS, since the randomness in SDE sampling often provides a highly diverse sample, which is not always less noisy, and could guide the 3D model in the wrong direction. Then motivated by the fact that an ordinary differential equation (ODE) of an SDE can provide a deterministic and consistent sampling trajectory, we propose a novel and effective “Consistent3D” by designing a consistency distillation sampling loss to distill the deterministic sampling prior into a 3D model for text-to-3D generation. Extensive experimental results show that our Consistent3D surpasses state-of-the-art methods in generating high-fidelity and diverse 3D objects and large-scale scenes, also achieving superior quantitative evaluation performance.
@@ B.2 Additional Results
More qualitative results for both the coarse and fine stages are shown in Figure 7. We notice that our Consistent3D can generate robust geometry in the coarse stage and then seemly enhance high-frequency and sophisticated details in the fine stage. Please check out the video demos for more generated results.
@@ Appendix C Theoretical Proof
[…]

### v2

@@ Abstract
Score distillation sampling (SDS) and its variants have greatly boosted the development of text-to-3D generation, but are vulnerable to geometry collapse and poor textures yet. To solve this issue, we first deeply analyze the SDS and find that its distillation sampling process indeed corresponds to the trajectory sampling of a stochastic differential equation (SDE): SDS samples along an SDE trajectory to yield a less noisy sample which then serves as a guidance to optimize a 3D model. However, the randomness in SDE sampling often leads to a diverse and unpredictable sample which is not always less noisy, and thus is not a consistently correct guidance, explaining the vulnerability of SDS. Since for any SDE, there always exists an ordinary differential equation (ODE) whose trajectory sampling can deterministically and consistently converge to the desired target point as the SDE, we propose a novel and effective “Consistent3D" method that explores the ODE deterministic sampling prior for text-to-3D generation. Specifically, at each training iteration, given a rendered image by a 3D model, we first estimate its desired 3D score function by a pre-trained 2D diffusion model, and build an ODE for trajectory sampling. Next, we design a consistency distillation sampling loss which samples along the ODE trajectory to generate two adjacent samples and uses the less noisy sample to guide another more noisy one for distilling the deterministic prior into the 3D model. Experimental results show the efficacy of our Consistent3D in generating high-fidelity and diverse 3D objects and large-scale scenes, as shown in Fig. 1. The codes are available at https://github.com/sail-sg/Consistent3D.
@@
footnotetext: Corresponding author.
@@ 1 Introduction
To address this critical issue, in this paper, we propose a novel and effective method, dubbed“Consistent3D”, which guides text-to-3D generation using deterministic sampling prior. As illustrated in Fig. 2, for these unpredictable and uncontrollable trajectories sampled from SDE solution distribution, there theoretically always exists a corresponding ordinary differential equation (ODE) whose trajectory shares the same marginal distributions with the SDE solution [41]. Importantly, this ODE trajectory is deterministic and consistently converges to the same target point as the SDE. Sampling along the ODE trajectory guarantees a predictable and deterministic next point, which always directs towards the desired target and thus providing a reliable and consistent guidance. This motivates us to explore the text-to-3D generation from the ODE deterministic sampling perspective.
@@ 2 Related Works
Text-to-3D Generation stands for generating 3D contents from a given text description. Current 3D generative models [29, 15], usually work in a single object category and suffer from limited diversity due to the lack of large-scale 3D datasets. To achieve open-vocabulary 3D generation, pioneered by DreamFusion [31], several approaches propose to lift text-image diffusion models [34] for 3D generation [56, 47, 57]. The key mechanism of such approaches is the score distillation sampling (SDS), where diffusion priors are used to supervise the optimization of a 3D representation. The following works continue to further improve the stability and fidelity of generation of various aspects, e.g., advanced 3D representation [3, 43, 51, 46], coarse-to-fine training strategy [20, 57, 47, 53] and 3D-aware diffusion priors [56, 38, 19, 22, 44].
@@ 4 Method
Here we elaborate on our proposed Consistent3D for effective text-to-3D generation. In Sec. 4.1, we reveal the underlying mechanism of Score Distillation Sampling (SDS) which aims to approximate the SDE sampling process and motivates our proposed methods. Then in Sec. 4.2, we introduce Consistency Distillation Sampling (CDS), a loss designed to efficiently distill deterministic sampling priors for text-to-3D generation, accompanied with a theoretical justification regarding the error bound. Finally, we present how to use our CDS to build our text-to-3D generation framework, dubbed “Consistent3D", in Sec. 4.3.
@@
       $t_{2}\leftarrow t_{\max}-(t_{\max}-t_{\min})\sqrt{{i}/{N}}$
       Sample camera pose $\pi$
      Sample $t_{1}\in{\mathcal{U}}[t_{2}+\delta,t_{2}+\Delta]$
      $\mathbf{x}_{\pi}\leftarrow g(\bm{\theta},\pi)$
      $\mathbf{x}_{t_{1}}\leftarrow\mathbf{x}_{\pi}+\sigma_{t_{1}}\bm{\epsilon}^{*}$
      $\mathbf{d}_{i}\leftarrow\left(\mathbf{x}_{t_{1}}-D_{\phi}(\mathbf{x}_{t_{1}},t_{1},y)\right)/\sigma_{t_{1}}$
      $\hat{\mathbf{x}}_{t_{2}}\leftarrow\mathbf{x}_{t_{1}}+(\sigma_{t_{2}}-\sigma_{t_{1}})\mathbf{d}_{i}$
      $\hat{\mathbf{x}}_{0}\leftarrow\mathbf{x}_{\pi}+\operatorname{sg}\left[\sigma_{t_{1}}(\bm{\epsilon}^{*}-\mathbf{d}_{i})\right]$
      ${\mathcal{L}}_{\text{CDS}}(\bm{\theta};\pi)\leftarrow\lambda(t_{2})\lVert\hat{\mathbf{x}}_{0}-\operatorname{sg}\left[D_{\phi}(\hat{\mathbf{x}}_{t_{2}},t_{2},y)\right]\rVert_{2}^{2}$
      $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}{\mathcal{L}}_{\text{CDS}}(\bm{\theta};\pi)$
@@ 4.1 Revisit Score Distillation Sampling
First, we discretize the reverse SDE in Eq. (2) and perform the stochastic sampling process following Ho et al. [11], which results in the SDE solution trajectory defined as
@@ 4.2 Consistency Distillation Sampling
Specifically, we define the time step schedule [53] of $t_{2}$ according to the current training iteration:
@@
Assume that the diffusion model $D_{\phi}(\cdot)$ satisfies the Lipschitz condition. Define $\Delta:=\sup|t_{1}-t_{2}|$. For any given camera pose $\pi$, if convergence is achieved according to Eq. (10), then there exists a corresponding real image $\mathbf{x}^{*}\sim p_{\text{data}}(\mathbf{x})$ such that
@@ 4.3 Consistent3D
NeRF Optimization Stage. We adopt multi-resolution hash grids, Instant NGP [28] to parameterize the scene by density and color with MLPs, which improves training and rendering efficiency. We follow Magic3D on density bias initialization, camera and light augmentation. In addition, we use orientation loss [31] and 2D normal smooth loss [24]. At this stage, we render $64\times 64$ images and use our proposed CDS as guidance. We set $t_{\max}=0.7$, $t_{\min}=0.1$, $\delta=0.1$, and $\Delta=0.2$.
Mesh Refinement Stage. We convert the neural field into Signed Distance Field (SDF) by subtracting it with a fixed threshold and then optimizing a high-resolution DMTet [37]. We also initialize the volume texture field directly with the color field from the coarse stage. In addition, we use normal consistency loss and Laplacian smoothness loss. In the refinement stage, we render $512\times 512$ images and set $t_{\max}=0.5$, $t_{\min}=0.02$, $\delta=0.1$, and $\Delta=0.1$.
@@ 5.1 Implementation Details
Consistent3D is implemented in PyTorch with a single NVIDIA A100 GPU based on threestudio [9] with Stable Diffusion v2.1 [34]. We use the Adan [49] optimizer with a learning rate of $0.05$ for grid encoder and $0.005$ for other parameters, and a weight decay of $2\times 10^{-8}$. Further implementation details are provided in Sec. B.1.
@@ 5.3 Comparison with the State-of-The-Art
Quantitative Results. In Tab. 1, we report the results of CLIP R-Precision [32] for 3D objects generated using $40$ randomly selected text prompts from the DreamFusion gallery. See more details in Sec. B.1. Each 3D object is rendered from $120$ viewpoints with a uniform azimuth angle. The CLIP R-Precision score is computed by averaging the similarity scores between each rendered view and the corresponding text prompt [52]. Additionally, we also conduct a head-to-head user study in Tab. 3. Our quantitative analysis shows the superior performance of our method.
@@ 5.4 Ablation Study
We present an ablation study to evaluate the effects of various components in our approach in Fig. 6 and Tab. 2. We conduct experiments with the following configurations: (a) a random time step schedule in DreamFusion [31]; (b) a predetermined time step schedule from Consistency Distillation [42]; (c) varied random noise in each iteration; and (d) our proposed method incorporating all components. The results in Fig. 6(a) reveal that a random time step schedule detrimentally affects both geometry and texture modeling, since it disrupts established rules of sampling process. Fig. 6(b) suggests that a predetermined time-step schedule is suboptimal for optimization-based methods, since gradient descent does not ensure monotonic optimization progress. This implies that minor randomness helps to accommodate these variations. Fig. 6(c) shows that fixed noise aids in better convergence by providing a consistent perturbation in each iteration.
@@ 6 Conclusion
In this work, we first connect Score Distillation Sampling (SDS), a leading text-to-3D generation approach, with the solution trajectory sampling of a stochastic differential equation (SDE). This connection helps us to understand the vulnerability in SDS, since the randomness in SDE sampling often provides a highly diverse sample, which is not always less noisy, and could guide the 3D model in the wrong direction. Then motivated by the fact that an ordinary differential equation (ODE) of an SDE can provide a deterministic and consistent sampling trajectory, we propose a novel and effective “Consistent3D" by designing a consistency distillation sampling loss to distill the deterministic sampling prior into a 3D model for text-to-3D generation. Extensive experimental results show that our Consistent3D surpasses state-of-the-art methods in generating high-fidelity and diverse 3D objects and large-scale scenes.
Acknowledgement
Pan Zhou was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant.
@@ B.2 Additional Results
(a) (b) (c) (d) Ours
0.319 0.325 0.340 0.348
Table 2: CLIP R-precision of the ablation study. (a) random time step schedule; (b) predetermined time step schedule; (c) random noise in each iteration; (d) our proposed configuration.
DF M3D PD Ours
User Prefer. Rate ($\uparrow$) 15.0% 5.4% 15.0% 64.6%
GPT4-V Avg. Rank ($\downarrow$) 3.0 3.8 1.8 1.4
Table 3: User/AI study. DF: DreamFusion; M3D: Magic3D; PD: ProlificDreamer.
More qualitative results for both the coarse and fine stages are shown in Fig. 7. We notice that our Consistent3D can generate robust geometry in the coarse stage and then seemly enhance high-frequency and sophisticated details in the fine stage. We conduct the ablation study quantitatively in Tab. 2, the results also support the superiority of our Consistent3D over others.
We also conduct a user study in Tab. 3 where $50$ users select best from multiple choices on $30$ generated 3D assets. Moreover, GPT4-V is used to evaluate and rank the generated 3D assets from aesthetic appeal, multi-view consistency and alignment with text prompt. Tab. 3 demonstrates the superiority of the results generated by our Consistent3D from both human and large multi-modal AI perspectives.
@@ Appendix C Theoretical Proof
Assume that the diffusion model $D_{\phi}(\cdot)$ satisfies the Lipschitz condition. Define $\Delta:=\sup|t_{1}-t_{2}|$. For any given camera pose $\pi$, if convergence is achieved according to Eq. (10), then there exists a corresponding real image $\mathbf{x}^{*}\sim p_{\text{data}}(\mathbf{x})$ such that